OpenStudy (anonymous):

In which function is x = 2 mapped to 32? g(x) =4(x+3)^2-68 h(x)=3x j(x)=2x-62

OpenStudy (campbell_st):

just substitute x = 2 into each equation. then the answer is the one that has a function value of 32. so g(2) = 4(2 + 3)^2 - 68 h(x) = 3 * 2 j(x) = 2*2 - 62

OpenStudy (anonymous):

which one is it?

OpenStudy (campbell_st):

well just calculate them out... it should be obvious which is isn't

OpenStudy (anonymous):

the third one right?

OpenStudy (campbell_st):

so 2 times 2 - 62 = 32...?

OpenStudy (campbell_st):

or... 4 - 62 = 32....

OpenStudy (anonymous):

ill do the next one just what is this one

OpenStudy (anonymous):

idk how to solve the fisrt equation

OpenStudy (campbell_st):

well use pedmas.... brackets 2 + 3 then square that number then multiply it by 4... then subtract 68